OpenStudy (anonymous):

Find the exact distance between the two given points. Simplify, if possible. Please show all of your work. (1,5) and (5,-3)

OpenStudy (anonymous):

\[Distance = \sqrt{(x2 - x1)^2 + (y2 - y1)^2}\]

OpenStudy (anonymous):

x1 = 1 x2 = 5 y1 = 5 y2 = -3

OpenStudy (anonymous):

Just plug in the numbers and simplify

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so it would be sqrt 80

OpenStudy (anonymous):

i dunno lemme check

OpenStudy (anonymous):

\[\sqrt{80} \space or \space 4\sqrt{5}\]

OpenStudy (anonymous):

I think you want the simplified one so i would write down 4sqrt{5}

OpenStudy (anonymous):

thanks you helped alot

OpenStudy (anonymous):

yw